OpenStudy (anonymous):

Write in logarithmic equation in exponential form. ln x=7 I think it's x=ln7

OpenStudy (jdoe0001):

http://www.purplemath.com/modules/logs/relatani.gif keep in mind \(\bf ln = log_e\)

OpenStudy (mathstudent55):

By definition of logarithm, \( \log_b{x} = y \) means \(b^y = x\) \( \ln x = 7\) \(\ln x\) simply means \( \log_e{x} \)

OpenStudy (mathstudent55):